# Xingjian Wu, Jianxin Jin, Wanghui Qiu, Peng Chen, Yang Shu, Bin Yang, Chenjuan Guo

### v2

@@ Appendix C Related Works
Time Series Analysis assumes a position of utmost significance within a wide array of fields, including the economy (Qiu et al., 2025b; a; Liu et al., 2025a), transportation (Wu et al., 2025c; 2024), health (Lu et al., 2023; 2024a; 2024b), weather (Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses numerous crucial tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024d), anomaly detection (Wang et al., 2025a; Qiu et al., 2025c; Wu et al., 2025b), classification (Liu et al., 2024d), imputation (Wu et al., 2023), and others (Qiu et al., 2025e; Wu et al., 2025a). Among these tasks, Time Series Forecasting is the one most widely applied in real-world situations.
@@
Time series forecasting (TSF) entails the prediction of future observations based on historical ones. Research findings have indicated that features learned through certain methods may exhibit superior performance compared to human-designed features (Dai et al., 2024c; Hu et al., 2025a; Dai et al., 2024d; Liu et al., 2025c; Sun et al., 2025b; a; Niu et al., 2025). By harnessing the representation learning capabilities of deep neural networks (DNNs), a multitude of deep-learning based approaches have emerged. For instance, TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, employing CNNs or RNNs to capture temporal dependencies. Moreover, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025c), LightGTS (Wang et al., 2025b), and MagicScaler (Pan et al., 2023), are capable of more accurately capturing the complex relationships between time points, thereby significantly enhancing forecasting performance. MLP-based methods, such as DUET (Qiu et al., 2025d), AMD (Hu et al., 2025a), SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters yet still manage to achieve highly competitive forecasting accuracy.

### v3

@@ Appendix E Related Works
Time Series Analysis assumes a position of utmost significance within a wide array of fields, including the economy (Qiu et al., 2025b; a; Liu et al., 2025a), transportation (Wu et al., 2025f; 2024), health (Lu et al., 2023; 2024a; 2024b), weather (Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses numerous crucial tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024c), anomaly detection (Wang et al., 2025a; Qiu et al., 2025c; Wu et al., 2025e), classification (Liu et al., 2024d), imputation (Wu et al., 2023; 2025b), and others (Qiu et al., 2025e; Liu et al., 2026; 2025c). Among these tasks, Time Series Forecasting is the one most widely applied in real-world situations.
@@
Time series forecasting (TSF) entails the prediction of future observations based on historical ones. Research findings have indicated that features learned through certain methods may exhibit superior performance compared to human-designed features (Hu et al., 2025a; Dai et al., 2024c; Liu et al., 2025d; Sun et al., 2025b; a; Niu et al., 2025). By harnessing the representation learning capabilities of deep neural networks (DNNs), a multitude of deep-learning based approaches have emerged. For instance, TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, employing CNNs or RNNs to capture temporal dependencies. Moreover, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025c), LightGTS (Wang et al., 2025b), and FLAME (Wu et al., 2025a), are capable of more accurately capturing the complex relationships between time points, thereby significantly enhancing forecasting performance. MLP-based methods, such as DUET (Qiu et al., 2025d), $K^{2}$VAE (Wu et al., 2025d), SRSNet (Wu et al., 2025c), AMD (Hu et al., 2025a), SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters yet still manage to achieve highly competitive forecasting accuracy.